OpenStudy (anonymous):

(2x-1)^0.5 + (x-1)^0.5 = 5 By square rooting them twice, i can get the x = 145, 5 is the root. But when i put 145 into the equation, it turn out not to be 145(but 5 really is the root) So how would i know which root should i choose, i mean how to prove the 145 isn't the root? Is it by just putting 145 into the equation? And when i use synthetic division on that equation with 145, i get the solution 0 ?? Does this mean synthetic division isn't applicable for some equations?

OpenStudy (anonymous):

ok first square both sides

OpenStudy (anonymous):

i get x = 145 , 5 in the end but 145 isn't the root

OpenStudy (anonymous):

how did u get that ?

OpenStudy (anonymous):

it must 5

OpenStudy (anonymous):

u should get 3x-2+2sqrt((x-1)(2x-1))=25

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i get 0 = x^2 - 150x + 725

OpenStudy (anonymous):

\[2\sqrt{(x-1)(2x-1)}=27-3x\]

OpenStudy (anonymous):

square this again

OpenStudy (anonymous):

ok I see there are two solution

OpenStudy (anonymous):

145 and 5

OpenStudy (anonymous):

u have to reject 145

OpenStudy (anonymous):

so we just show them when we put 145 in the equation the result doesn't make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok